### v2

@@ 6 The Dutch-book theorem for complete consistency
$\mathbf{3\Rightarrow 1)}$
The proof that an agent with two different GORs between a state $s$ and a state $s^{\prime}$ can be Dutch-booked resembles the classical Dutch-book arguments of De Finetti and Ramsey. Therefore, the proof is deferred to the Appendix. $\blacksquare$
The next example illustrates the theorem.
Example 5
@@
This is a variation of the famous Monty Hall problem, in which the initially selected door is determined at random and not revealed to the decision maker. There are three doors. A car is placed behind one of the doors, and a goat is placed behind each of the other two doors. The state represents the door with the car: $\mathcal{S}=\left\{A,B,C\right\}$. Then, one door, call it $d$, is drawn from a uniform distribution (in the Monty Hall problem, this is the door assigned to the agent). Finally, another door behind which there is a goat, call it $h\not=d$, is opened; it is the first non-car door preceding $d$ in an $a$-$b$-$c$ circle. So, for instance, conditional on $s=B$, door $a$ is opened with probability $2/3$ (the probability that $d\in\left\{b,c\right\}$), and door $c$ is opened with probability $1/3$ (the probability that $d=a$). The agent only observes which door is opened, so the possible contingencies are: =$\mathcal{H}=\left\{a,b,c\right\}$. Overall, conditional on each $s\in\mathcal{S}$, the probabilities of each $h\in\mathcal{H}$ are:
@@
$s=A$ 00 $\frac{1}{3}$ $\frac{2}{3}$
$s=B$ $\frac{2}{3}$ 00 $\frac{1}{3}$
$s=C$ $\frac{1}{3}$ $\frac{2}{3}$ 00
@@
First, suppose that, at every $h\in\mathcal{H}$, the agent assigns equal probability to each of the two states in $S(h)$. Then, the agent can be Dutch-booked (in objective expected terms). To see this, compute the three discounted odds ratios:
$o(B,C|a)$ $=$ $\frac{\Pr(a|C)}{\Pr(a|B)}=\frac{1/3}{2/3}=\frac{1}{2};$
$o(C,A|b)$ $=$ $\frac{\Pr(b|A)}{\Pr(b|C)}=\frac{1/3}{2/3}=\frac{1}{2};$
$o(B,A|c)$ $=$ $\frac{\Pr(c|A)}{\Pr(c|B)}=\frac{2/3}{1/3}=2.$
@@
$g(C|a)$ $=$ 3, $3,\text{ \ }g(B|a)=-2;$
$g(A|b)$ $=$ 3, $3,\text{ \ }g(C|b)=-2;$
$g(B|c)$ $=$ 3, $2\displaystyle 3,\text{ \ }g(A|c)=-$ 2 .
@@
Second, suppose that the agent derives his beliefs as follows. He has a uniform prior, and, thinking of the correct updating in the original Monty Hall problem, believes that the car is more likely to be behind the unopened door that has not been assigned to him (i.e., not door $d$). For instance, suppose that $h=a$ and the agent correctly computes
$\Pr\left(d=b|h=a\right)$
$=$ $\frac{\Pr(h=a|d=b)\Pr(d=b)}{\Pr(h=a|d=b)\Pr(d=b)+\Pr(h=a|d=c)\Pr(d=c)}$
$=$ $\frac{\Pr(h=a|d=b,s=B)+\Pr(h=a|d=b,s=C)}{\left(\begin{array}[]{c}\Pr(h=a|d=b,s=B)+\Pr(h=a|d=b,s=C)\\ +\Pr(h=a|d=c,s=B)+\Pr(h=a|d=c,s=C)\end{array}\right)}$
$=$ $\frac{2}{2+1}=\frac{2}{3}.$
So, he has $\mu(s=B|h=a)<1/2$. However, given $h=a$, the car is more likely to be behind door $b$:
$\Pr\left(s=B|h=a\right)$
$=$ $\frac{\Pr(h=a|s=B)\Pr(s=B)}{\Pr(h=a|s=B)\Pr(s=B)+\Pr(h=a|s=C)\Pr(s=C)}$
$=$ $\frac{\Pr(s=B)}{\Pr(s=B)+\Pr(d=b)\Pr(s=C)}$
$=$ $\frac{1}{1+\frac{1}{3}}=\frac{3}{4}.$
In other words, the agent perceives the opposite correlation between the fundamental state and the observed contingency. This increases the discrepancy between generalized odds ratios, making it easier for the bookmaker to Dutch-book the agent. [^13] To be precise, without limited liability any discrepancy allows the bookmaker to extract an infinite amount of money (in expectation) from the agent. But under limited liability, the higher discrepancy allows to strictly increase the bookmaker’s expected profit.
@@ 7 Conclusion
Dutch-book arguments have already been used to motivate probabilistic sophistication and conditioning. We introduce novel Dutch-book arguments that motivate two forms of belief consistency in a learning framework a la Anscombe-Aumann, with separation between subjective uncertainty (the state) and objective uncertainty (lotteries over learning paths conditional on the state). Bayes-plausibility is necessary and sufficient to avoid being Dutch-booked in a deterministic sense, i.e., under each realized state and learning path. A stronger form of coherence, guaranteed by deriving all beliefs from a subjective lexicographic prior with the correct models, is necessary and sufficient to avoid being Dutch-booked “in objective expected terms”, i.e., under each realized state, given the corresponding lottery on the learning paths. The latter form of Dutch-booking entails sure aggregate losses across a population of agents, thus showing the perils of heterogenous priors across different individuals, and of model misspecification even in absence of an objective prior.
@@ Appendix A Appendix
In the context of a sequential game with complete information and perfect recall, Battigalli et al. (2023) introduce a classification of the belief systems of a player over the opponents’ strategies. We study their classification of belief systems as a special case of the analysis of the main body of the paper. In a game, the moves of a player are independent of the moves of the co-players that are yet to be observed, and the two jointly determine the path of the game, thus the learning path. Therefore, the assumption that every contingency is reached with the same probability under every consistent state is satisfied in that context. For concreteness, we stick to the sequential games application, but the analysis extends seamlessly to situations where this assumption is satisfied, for instance in situations where the an agent does not choose what to observe but the factors that determine the observations are independent of the state itself, as in the running example of the paper.
Let $S_{-i}$ be the set of stategy profiles of the opponents of player $i$. Let $H_{i}$ be the arborescence of information sets of player $i$, endowed with the precedence relation $\prec$. Each information set $h_{i}$ corresponds to the observation of some past moves of the opponents, summarized by the event $S_{-i}(h_{i})\subseteq S_{-i}$. A belief system is an array $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ of probability measures over $S_{-i}$ such that $\mu_{i}(S_{-i}(h_{i})|h_{i})=1$ for every $h_{i}\in H_{i}$.
Taking as primitive a collection of conditioning events $\mathcal{C\subseteq}2^{S_{-i}}\backslash\left\{\emptyset\right\}$, a conditional probability system over $S_{-i}$ is an array of probability measures $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in\mathcal{C}}$ such that $\bar{\mu}_{i}(C|C)=1$ for every $C\in\mathcal{C}$, and for every $D\in\mathcal{C}$ with $D\subset C$, for each $E\subseteq D$,
@@
A conditional probability system is complete when it is defined with respect to the collection $\mathcal{C=}2^{S_{-i}}\backslash\left\{\emptyset\right\}$.
@@
[…]
$\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i})).$
The definition of complete consistency of Battigalli et al. (2023) uses CCPS’s to discipline beliefs at counterfactual contingencies. An information set $h$ does not carry any additional information about the strategies of the opponents besides ruling out all the strategies that do not belong to $S_{-i}(h)$; for this reason, it is appropriate to derive each belief by mere conditioning in place of Bayes rule, from the most general theory that explains the contingency.
@@
For every information set $h_{i}$, for every pair $s_{-i},s_{-i}^{\prime}\in S_{-i}(h_{i})$, and for every probability distribution over player $i$’s strategies, the probability of reaching the information set is identical under $s_{-i}$ and $s_{-i}^{\prime}$: $p(h_{i}|s_{-i})=p(h_{i}|s_{-i}^{\prime})$.
@@
Although the use of CCPSs is appropriate in this context, one might as well model the theories from which a player derives just as a LCPS, and derive each belief by conditioning the first theory in the list that is consistent with the contingency; the two definitions of complete consistency are equivalent. [^14] The equivalence between LCPSs and CCPSs is well-known. We formalize and prove this result for the analysis to be self-contained.
@@
[…]
    $\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i}));$
@@
    there exists a LCPS $\hat{\mu}_{i}=(\mu_{i}^{1},...,\mu_{i}^{n})$ such that for each $h_{i}\in H_{i}$ and $s_{-i}\in S_{-i}(h_{i})$,
    $\mu_{i}(s_{-i}|h_{i})=\frac{\mu_{i}^{k}(s_{-i})}{\mu_{i}^{k}(S_{-i}(h_{i}))},$

### v3

@@ 6 The Dutch-book theorem for complete consistency
$\mathbf{3\Rightarrow 1)}$ It will follow as an implication of the more general Proposition 1. $\blacksquare$
6.1 Robustness to the Bookmaker Misspecification
One natural concern about the relevance of the previous result is that a misspecified agent has been shown to be exploitable by a bookmaker who perfectly knows the true data-generating process. In principle, if the bookmaker model is also incorrect, then the proposed bet may induce some losses for the bookmaker in some $s\in S$. We next show that this is not the case as long as the misspecification is sufficiently small, and we quantify what small means. To do so, we start by observing that beliefs that are not completely coherent admit a special form of inconsistency in the generalized discounted odd ratios.
Lemma 1
If $\mu$ is not completely coherent, then there exists a state $s$ and a GOR between $s$ and itself with value $r<1$ and where $s^{0},...,s^{n-1}$ are all distinct for each $k\in\left\{i,...,n-1\right\}$, $h^{k}\not=h^{k+1}$.
The lemma shows that without complete coherence, the agent’s belief can be used to build a GOR that starts and ends at the same state and differs from $1$, the only value it could instead assume under complete coherence. This observation motivates the following definition.
Definition 10
Let $r<1$ and $n\geq 2$. We say that $\mu$ is $(r,n)$-not-coherent if there exists a state $s$ and a GOR between $s$ and itself with value $r<1$ and where $s^{0},...,s^{n-1}$ are all distinct for each $k\in\left\{i,...,n-1\right\}$, $h^{k}\not=h^{k+1}$.
This definition quantifies the degree of incoherence between the agent’s beliefs $\mu$ and the data generating process $p$, with smaller values of $r$ and $n$ indicating a more significant incoherence. Indeed, a smaller $r$ indicates that a larger departure from $1$ of the GOR between $s$ and itself, and a smaller $n$ indicates that such a departure can be obtained by just using a smaller number of comparisons between state-contingency pairs. Next, we define what it means for a Dutch book to be robust to misspecification on the bookmaker side.
Definition 11
Let $\beta\in\mathbb{R}_{+}$. A Dutch book $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a $\beta$-robust Dutch book if for every $\hat{p}$ such that
$\frac{\hat{p}(h|s)}{\hat{p}(h^{\prime}|s)}\frac{p(h^{\prime}|s)}{p\left(h|s\right)}<\beta\qquad\forall h,h^{\prime}\in\mathcal{H},s\in S$
we have that for all $s\in\mathcal{S}$,
$\sum_{h\in\mathcal{H}}\hat{p}(h|s)g(s|h)\leq 0,$
and for some $s\in\mathcal{S}$,
$\sum_{h\in\mathcal{H}}\hat{p}(h|s)g(s|h)<0.$
This refinement of the Dutch book concept should be interpreted as follows. As before, the agent is endowed with beliefs $\mu$, and the bookmaker considers the data generating process $p$. However, she no longer completely trusts it and envisions the possibility that their understanding of how a state translates into different consequences (i.e., $\frac{p(h^{\prime}|s)}{p\left(h|s\right)}$) may be incorrect to some extent. Therefore, they require the Dutch book to lead to a payoff that is nonnegative under every state and positive under some state, even if the true data generating process is some different $\hat{p}$ that is sufficiently close to $p$.
The next result shows that such robustness can be achieved, and it links it to the quantitative notion of incoherence defined above.
Proposition 1
Suppose that $\mu$ is $(r,n)$-not-coherent. Then for every $\delta>0$ there exists a $\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$-robust Dutch book.
Importantly, the result shows that to achieve a Dutch book, what is key is the comparison between the disagreement between the agent and the bookmaker, and the misspecification of the bookmaker. If the disagreement between the two, captured by the measure $\left(r,n\right)$ of how far are beliefs $\mu$ from being generated by the bookmaker model $p$ and some prior, is sufficiently high compared to the extent of misspecification of $p$ with respect to $\hat{p}$.
7 Applications to Correlation Neglect
We next apply our main theorem to one of the most pervasive and studied forms of misspecification, correlation neglect. We start by studying how the use of different financial instruments in venture capital can be viewed as a Dutch book due to neglect of correlation on the startup side. We then move to see how some offers in a variant of the classical Monty Hall problem can also be interpreted in the same way.
Financial Instruments
Consider a binary-state environment, where states $\mathcal{S}=\left\{G,B\right\}$ indicate whether a startup is good or bad. Evidence about the startup’s quality accumulates over time through a sequence of binary signals indicating whether key milestones are achieved. That is, the set of contingencies is $\mathcal{H=}\left\{\left\{\gamma,\beta\right\}^{n}\right\}_{n\in\left\{1,...,N\right\}}$, $N\geq 4$, where $\gamma$ was a good performance and $\beta$ was a bad performance in a key milestone. The objective data generating process over contingencies is Markov, with two key properties. First, good firms are more likely to achieve the milestones, and second, there is persistence in the milestone process. Regardless of whether the startup is good or bad, the probability of achieving the second milestone is higher if the first one has been achieved. For concreteness, the objective DGP is recursively given by
$p\left(\gamma|G\right)$ $=$ $\frac{2}{3}=p\left(\beta|B\right)$
$p\left(\left(h,\gamma,\gamma\right)|\left(h,\gamma\right),G\right)$ $=$ $\frac{3}{4}=p\left(\left(h,\beta,\beta\right)|\left(h,\beta\right),B\right)\quad\forall h\in\left\{\gamma,\beta\right\}^{n},n\leq N-2$
$p\left(\left(h,\beta,\gamma\right)|\left(h,\beta\right),G\right)$ $=$ $\frac{1}{2}=p\left(\left(h,\gamma,\beta\right)|\left(h,\gamma\right),B\right)\quad\forall h\in\left\{\gamma,\beta\right\}^{n},n\leq N-2\text{.}$
The startup updates its beliefs using Bayes’ rule and understands that milestones are more likely to be reached when the state is good than when it is bad. However, the startup suffers from correlation neglect and treats milestone achievements as i.i.d. conditional on the state. Concretely, the belief system of the startup is given by
$\mu\left(G|\varnothing\right)=\frac{1}{2}\text{,}$
and for all $h\in\left\{\gamma,\beta\right\}^{n},n\leq N$
$\mu\left(G|h\right)=\frac{\left(\frac{2}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\gamma\right\}|}\left(\frac{1}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\beta\right\}|}}{\left(\frac{2}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\gamma\right\}|}\left(\frac{1}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\beta\right\}|}+\left(\frac{2}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\beta\right\}|}\left(\frac{1}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\lambda\right\}|}}\text{.}$ (10)
By Theorem 1, the startup cannot be deterministically Dutch booked, as the beliefs are derived using Bayesian updating from the conditional probabilities
$\widetilde{p}\left(\left(h,\gamma\right)|h,G\right)$ $=$ $\frac{2}{3}\qquad\forall h\in\left\{\gamma,\beta\right\}^{n},n\leq N-1$
$\widetilde{p}\left(\left(h,\gamma\right)|h,B\right)$ $=$ $\frac{1}{3}\qquad\forall h\in\left\{\gamma,\beta\right\}^{n},n\leq N-1\text{.}$
However, a venture capital offering a combination of “purchase the firm” and “convertible debt” can Dutch-book the startup by exploiting their correlation neglect. In particular, and without loss of generality, given the risk neutrality of the agent, suppose that ownership of a good firm is worth $\frac{2}{3}$ in the good state and $\frac{1}{3}$ in the bad state (that is, each state is worth its limit distribution over positive performances). The startup initially holds ownership. The venture capitalist offers two different contracts in contingencies of length $4$. After the sequence $\left(\gamma,\beta,\beta,\gamma\right)$, the venture capitalist offer to buy the startup at a price of $0.51$. This corresponds to the bet
$g\left(G|\left(\gamma,\beta,\beta,\gamma\right)\right)=0.51-\frac{2}{3}\text{ and }g\left(B|\left(\gamma,\beta,\beta,\gamma\right)\right)=0.51-\frac{1}{3}\text{.}$
By equation (10), $\mu\left(G|\left(\gamma,\beta,\beta,\gamma\right)\right)=\frac{1}{2}$, and thus the startup accepts and sells ownership, overweighting the informativeness of the two consecutive failures.
Instead, after sequence $\left(\gamma,\gamma,\gamma,\gamma\right)$, convertible debt at a rate below the market rate is offered to the up-and-coming startup. This below-market-rate debt amounts to a state-independent financing advantage of $\frac{1}{100}$ to the startup, but in the unlikely case the startup turns out to be of the bad type, the venture capital gains control of half of the ownership of the firm. This corresponds to the bet [^16] This contract can also be approximately implemented as a transfer of property if the fraction of milestones over a long horizon is smaller than $1/2$, as this is an event that has probability approaching $1$ under $B$ and 0 under $G$ as the horizon grows.
$g\left(G|\left(\gamma,\gamma,\gamma,\gamma\right)\right)=\frac{1}{100}\text{ and }g\left(B|\left(\gamma,\gamma,\gamma,\gamma\right)\right)=\frac{1}{100}-\frac{1}{2}\frac{1}{3}\text{.}$
By equation (10), $\mu\left(G|\left(\gamma,\gamma,\gamma,\gamma\right)\right)=\frac{16}{17}$ and thus the startup accepts the convertible debt, overweighting the informativeness of these successes. Easy computations verify that this is indeed a Dutch book. [^17] In state $G$, contingency $\left(\gamma,\beta,\beta,\gamma\right)$ is reached with probability $0.0416$ with a startup payoff of 0.1566-0.1566, and $\left(\gamma,\gamma,\gamma,\gamma\right)$ is reached with probability $0.28125$ and a startup payoff of $0.01$. In state $B$, contingency $\left(\gamma,\beta,\beta,\gamma\right)$ is reached with probability $0.03125$ with a startup payoff of $0.1766$, and $\left(\gamma,\gamma,\gamma,\gamma\right)$ is reached with probability $0.0416$ and a startup payoff of $0.01-\frac{1}{6}$, resulting in strictly negative expected losses in both states. No matter the baseline frequency of successful startups, the venture capitalist can, on average, extract positive returns from both types, exploiting the acquisition of good startups after mixed performance and convertible debt issued to bad startups after an initial sequence of successes.
We also observe that, consistently with Theorem 2, the generalized odds ratios of the startup do not agree
$o\left(G,B|\left(\gamma,\gamma,\gamma,\gamma\right)\right)$ $=$ $\frac{\mu\left(g|\left(\gamma,\gamma,\gamma,\gamma\right)\right)}{p\left(\left(\gamma,\gamma,\gamma,\gamma\right)|G\right)}\frac{p\left(\left(\gamma,\gamma,\gamma,\gamma\right)|B\right)}{\mu\left(b|\left(\gamma,\gamma,\gamma,\gamma\right)\right)}=\frac{16/17}{0.28125}\frac{0.0416}{1/17}=2.366$
$>$ $\frac{1/2}{0.0416}\frac{0.03125}{1/2}=\frac{\mu\left(G|\left(\gamma,\beta,\beta,\gamma\right)\right)}{p\left(\left(\gamma,\beta,\beta,\gamma\right)|G\right)}\frac{p\left(\left(\gamma,\beta,\beta,\gamma\right)|B\right)}{\mu\left(B|\left(\gamma,\beta,\beta,\gamma\right)\right)}=o\left(G,B|\left(\gamma,\beta,\beta,\gamma\right)\right)\text{.}$
In words, the agent interprets the good type sequence $\left(\gamma,\gamma,\gamma,\gamma\right)$ too positively, as it fails to account for the persistence of performance.
This application, among other things, illustrates the difference between our Dutch book argument and preexisting arguments against misspecification. Typically, it is argued that some misspecifications cannot persist because, in the long run, key statistics of the data force the agent to either reject or incompletely trust their model. However, the form of misspecification above would be immune to these arguments. Indeed, as the horizon grows, i.e., $N\rightarrow\infty$, the startup beliefs concentrate on the actual state, i.e., $\mu\left(G|h\right)$ converges to $1$ almost surely conditional on being in state $G$, and the same holds for $\mu\left(B|h\right)$ under $B$. Moreover, the startup will observe a fraction of success equal to the one predicted by their incorrect model (i.e., the stationary distribution of $\gamma$ under $g$ is $2/3$, which is the probability of $\gamma$ under their i.i.d. model). Under essentially all the work that studied the persistence of misspecification, these two conditions are sufficient to guarantee that their misspecified model will not be relaxed or changed (Gagnon-Bartsch, Rabin, and Schwartzstein 2023, Fudenberg and Lanzani 2023, Ba 2025, He and Libgober 2025, Lanzani 2025). Our argument against misspecification, instead, does not work at the level of beliefs and how they evolve over time, but at the level of material losses deriving from the cross-sectional exploitation of a population of misspecified agents.
Monty Hall offers
@@ 7 Applications to Correlation Neglect
This is a variation of the famous Monty Hall problem, in which the initially selected door is determined at random and not revealed to the decision maker. There are three doors. A car is placed behind one of the doors, and a goat is placed behind each of the other two doors. The state represents the door with the car: $\mathcal{S}=\left\{A,B,C\right\}$. Then, one door, call it $d$, is drawn from a uniform distribution (under the Monty Hall interpretation, this is the door initially assigned to the agent). Finally, another door behind which there is a goat, call it $h\not=d$, is opened; it is the first non-car door preceding $d$ in an $a$-$b$-$c$ cycle. So, for instance, conditional on the state being $B$, door $a$ is opened with probability $2/3$ (the probability that $d\in\left\{b,c\right\}$), and door $c$ is opened with probability $1/3$ (the probability that $d=a$). The agent only observes which door is opened, so the possible contingencies are: $\mathcal{H}=\left\{a,b,c\right\}$. Overall, conditional on each $s\in\mathcal{S}$, the probabilities of each $h\in\mathcal{H}$ are:
@@
$s=A$ 0 $\frac{1}{3}$ $\frac{2}{3}$
$s=B$ $\frac{2}{3}$ 0 $\frac{1}{3}$
$s=C$ $\frac{1}{3}$ $\frac{2}{3}$ 0
@@
First, suppose that, at every $h\in\mathcal{H}$, the agent assigns equal probability to each of the two states in $S(h)$. That is, they understand that the car cannot be behind the open door, but they neglect the correlation between the open door and the one containing the car, induced by the open-in-cycle mechanism. Then, the agent can be Dutch-booked (in objective expected terms). To see this, compute the three discounted odds ratios:
$o(B,C|a)$ $=$ $\frac{p(a|C)}{p(a|B)}=\frac{1/3}{2/3}=\frac{1}{2};$
$o(C,A|b)$ $=$ $\frac{p(b|A)}{p(b|C)}=\frac{1/3}{2/3}=\frac{1}{2};$
$o(B,A|c)$ $=$ $\frac{p(c|A)}{p(c|B)}=\frac{2/3}{1/3}=2.$
@@
$g(C|a)$ $=$ $3,\text{ \ }g(B|a)=-2;$
$g(A|b)$ $=$ $3,\text{ \ }g(C|b)=-2;$
$g(B|c)$ $=$ $3,\text{ \ }g(A|c)=-2.$
@@
Second, suppose that the agent derives his beliefs as follows. He has a uniform prior, and, thinking of the correct updating in the original Monty Hall problem, believes that the car is more likely to be behind the unopened door that has not been assigned to him (i.e., not door $d$). For instance, suppose that we are in contingency $a$ and the agent correctly computes
$\Pr\left(d=b|h=a\right)$ $=$ $\frac{\Pr(h=a|d=b)\Pr(d=b)}{\Pr(h=a|d=b)\Pr(d=b)+\Pr(h=a|d=c)\Pr(d=c)}$
$=$ $\frac{\Pr(h=a|\left(d,s\right)=\left(b,B\right))+\Pr(h=a|\left(d,s\right)=\left(b,C\right))}{\left(\begin{array}[]{c}\Pr(h=a|\left(d,s\right)=\left(b,B\right))+\Pr(h=a|\left(d,s\right)=\left(b,C\right))\\ +\Pr(h=a|\left(d,s\right)=\left(c,B\right))+\Pr(h=a|\left(d,s\right)=\left(c,C\right))\end{array}\right)}=\frac{2}{2+1}=\frac{2}{3}.$
So, he has $\mu(s=B|h=a)<1/2$, as prescribed by correct reasoning in the (original) Monty Hall problem: the car is less likely to be behind the door that is more likely to have been assigned to him. However, given contingency $a$, if the distribution over states is uniform, the car is more likely to be behind door $b$:
$\Pr\left(s=B|h=a\right)$ $=$ $\frac{p\left(a|B\right)\Pr(s=B)}{p\left(a|B\right)\Pr(s=B)+p\left(a|C\right)\Pr(s=C)}$
$=$ $\frac{\frac{2}{3}\Pr(s=B)}{\frac{2}{3}\Pr(s=B)+\Pr(d=b)\Pr(s=C)}=\frac{\frac{2}{3}}{\frac{2}{3}+\frac{1}{3}}=\frac{2}{3}.$
In other words, the agent perceives the opposite correlation between the fundamental state and the observed contingency. This increases the discrepancy between generalized odds ratios, making it easier for the bookmaker to Dutch-book the agent. [^18] To be precise, without limited liability, any discrepancy allows the bookmaker to extract an infinite amount of money (in expectation) from the agent. However, under limited liability, the higher discrepancy allows for a strict increase in the bookmaker’s expected profit.
@@ 8 Conclusion
Dutch-book arguments have been used to motivate probabilistic sophistication and conditioning. We introduce novel Dutch-book arguments that motivate two forms of belief consistency in a learning framework a la Anscombe-Aumann, with separation between subjective uncertainty (the state) and objective uncertainty (lotteries over learning paths conditional on the state). Bayes-plausibility is necessary and sufficient to avoid being Dutch-booked in a deterministic sense, i.e., under each realized state and learning path. A stronger form of coherence, guaranteed by deriving all beliefs from a subjective lexicographic prior with the correct models, is necessary and sufficient to avoid being Dutch-booked “in objective expected terms”, i.e., under each realized state, given the corresponding lottery on the learning paths. The latter form of Dutch-booking entails sure aggregate losses across a population of agents, thereby showing the perils of heterogeneous priors across individuals and of model misspecification even in the absence of an objective prior.
@@ Appendix A Appendix
In the context of a sequential game with complete information and perfect recall, Battigalli et al. (2023) introduce a classification of the belief systems of a player over the opponents’ strategies. We study their classification of belief systems as a special case of the analysis of the main body of the paper. In a game, the moves of a player are independent of the moves of the co-players that are yet to be observed, and the two jointly determine the path of the game, thus the learning path. Therefore, the assumption that every contingency is reached with the same probability under every consistent state is satisfied in that context. [^19] For concreteness, we stick to the sequential games application, but the analysis extends seamlessly to situations where this assumption is satisfied, for instance in situations where the an agent does not choose what to observe but the factors that determine the observations are independent of the state itself, as in the running example of the paper.
Let $S_{-i}$ be the set of strategy profiles of the opponents of player $i$. Let $H_{i}$ be the arborescence of information sets of player $i$, endowed with the precedence relation $\prec$. Each information set $h_{i}$ is associated with the event $S_{-i}(h_{i})\subseteq S_{-i}$ of the opponents strategies that reach $h_{i}$, under some strategy of player $i$. A belief system is an array $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ of probability measures over $S_{-i}$ such that $\mu_{i}(S_{-i}(h_{i})|h_{i})=1$ for every $h_{i}\in H_{i}$.
[…]
@@
[…]
@@
A belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ over the opponents strategies is completely consistent if there exists a CCPS $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in 2^{S_{-i}}\backslash\left\{\emptyset\right\}}$ such that
$\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i}))\qquad\forall h_{i}\in H_{i}.$
The definition of complete consistency of Battigalli et al. (2023) uses CCPS’s to discipline beliefs in counterfactual contingencies. An information set $h$ does not carry any additional information about the strategies of the opponents besides ruling out all the strategies that do not belong to $S_{-i}(h)$; for this reason, it is appropriate to derive each belief by mere conditioning in place of Bayes rule, from the most general theory that explains the contingency.
@@
For every $h_{i}\in H_{i}$, for every pair $s_{-i},s_{-i}^{\prime}\in S_{-i}(h_{i})$, and for every probability distribution over player $i$’s strategies, the probability of reaching the information set is identical under $s_{-i}$ and $s_{-i}^{\prime}$: $p(h_{i}|s_{-i})=p(h_{i}|s_{-i}^{\prime})$.
@@
Although the use of CCPSs is appropriate in this context, one might as well model the theories from which a player derives beliefs as a LCPS, and derive each belief by conditioning the first theory in the list that is consistent with the contingency; the two definitions of complete consistency are equivalent. [^20] The equivalence between LCPSs and CCPSs is well-known. We formalize and prove this result for the analysis to be self-contained.
@@
    there exists a CCPS $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in 2^{S_{-i}}\backslash\left\{\emptyset\right\}}$ such that
    $\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i}))\qquad\forall h_{i}\in H_{i};$
@@
    there exists a LCPS $\hat{\mu}_{i}=(\mu_{i}^{1},...,\mu_{i}^{n})$ such that
    $\mu_{i}(s_{-i}|h_{i})=\frac{\mu_{i}^{k}(s_{-i})}{\mu_{i}^{k}(S_{-i}(h_{i}))}\qquad\forall h_{i}\in H_{i},\forall s_{-i}\in S_{-i}(h_{i}),$
@@
$\mathbf{2\Rightarrow 1)}$ For each $C\in 2^{S_{-i}}\backslash\left\{\emptyset\right\}$, derive $\bar{\mu}_{i}(\cdot|C)$ by conditioning the first measure in the LCPS that assigns positive probability to $C$, so that if $C=S_{-i}(h_{i})$ for some $h_{i}\in H_{i}$, then $\bar{\mu}_{i}(\cdot|C)=\mu_{i}(\cdot|h_{i})$. There remains to show that $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in 2^{S_{-i}}\backslash\left\{\emptyset\right\}}$ is a CCPS. Now fix $D\subset C$ such that $\bar{\mu}_{i}(D|C)>0$. Thus, if $\bar{\mu}_{i}(\cdot|C)$ is derived from $\mu_{i}^{k}$, $\mu_{i}^{k}(D)>0$, and $\mu_{i}^{j}(D)\leq\mu_{i}^{j}(C)=0$ for every $j<k$. Therefore, also $\bar{\mu}_{i}(\cdot|D)$ is derived from $\mu_{i}^{k}$ and the chain rule is satisfied.